### v2

# BCS-BEC crossover of the strongly interacting 6Li-40K mixture

Abstract
We present quantum Monte Carlo calculations of the properties of a two-component mass imbalanced Fermi gas, corresponding to the 6Li-40K mixture. We compute the equation of state of the unpolarized system as a function of the scattering length with particular attention paid to the unitary limit, where the effect of the effective range of the interaction is explored. In order to investigate differences from the equal-mass case we also compute the pair-distribution function and the momentum distribution over a range of interaction strengths, which can provide information about the structure of the system. Finally, we compute the heavy-light quasiparticle spectrum for the full crossover regime. Our theoretical predictions, based on quantum Monte Carlo calculations, should inform future theoretical studies and can be tested by experiments with ultracold fermionic gases.
preprint: APS/123-QED
I Introduction
In recent years much work has been done to study ultracold Fermi gases, both experimentally and theoretically (for a complete review see for example Ref. Giorgini et al. (2008)). Through the use of a Feshbach resonance it is possible to tune the interaction between fermions as desired Chin et al. (2010), and the amazing experimental control on these systems gives access to beautiful experiments in which Fermi gases can be studied under a range of conditions. It is possible to probe the entire BCS-BEC crossover Strinati et al. (2018), where the system can be tuned from a BCS state where fermions are weakly interacting, to the unitary regime where the two-body scattering length is infinite, to a Bose-Einstein condensate (BEC) of bosons. When the population imbalance changes, the system can eventually exhibit different phases Zwierlein et al. (2006); Partridge et al. (2006) where superfluid and normal phases may coexist or phase separate Pilati and Giorgini (2008); Lobo et al. (2006) or other intriguing more exotic phases like the Larkin–Ovchinnikov–Fulde–Ferrell (LOFF) phase or $p$-wave superfluidity could appear Bulgac et al. (2006); Bulgac and Forbes (2008).
Considerable experimental effort has been spent on the investigation of trapped mass imbalanced two-component Fermi gases. Early work explored mainly systems consisting of a mixture of 6Li and 40K atoms  Wille et al. (2008); Taglieber et al. (2008); Spiegelhalder et al. (2010); Tiecke et al. (2010); Voigt et al. (2009); Trenkwalder et al. (2011), and in recent years experimental probes have been conducted for a wide variety of mass-imbalanced systems Khramov et al. (2014); Ravensbergen et al. (2018, 2020); Neri et al. (2020); Ciamei et al. (2022); Green et al. (2020); Schäfer et al. (2023) with various mass ratios. There has also been continued interest in probes of the 6Li-40K mixture, such as investigations into dimer stability Jag et al. (2016).
The properties of the mass-imbalanced system can be very different from the equal-mass case; for example, with a majority light population the Chandrasekhar-Clogston limit is very small and close to zero Gezerlis et al. (2009). Then, at unitarity, a two-component Fermi gas with unequal masses may exhibit very different properties with respect to the equal-mass case Gubbels et al. (2009); Baarsma et al. (2010), even at the mean field level Baranov et al. (2008); Wang et al. (2017). Systems of mass-imbalanced fermions also show new features in the few-body sector. The binding energy of an impurity was calculated in Ref. Combescot et al. (2007); however, finite systems with few-heavy atoms that interact with a single light one through a two-body potential with infinite scattering length have very intriguing nonuniversal properties Blume and Daily (2010a, b); Castin et al. (2010), and this could effectively modify the many-body structure of the state with polarization, supporting new phases that do not show in the equal mass case Braun et al. (2014); Roscher et al. (2015); Wang et al. (2017); Pini et al. (2021). It has also been shown that for very large mass ratios the ground state of the unpolarized system might be solid Petrov et al. (2007); Osychenko et al. (2012). In addition, there is rich physics to explore in mass-imbalanced systems existing in lower dimensions, such as novel types of pairing Rammelmüller et al. (2020) and quarteting Liu et al. (2023).
Since experiments are possible at very low temperatures, of the order of fractions of $T_{F}$, the zero-temperature effects can directly be compared to or extrapolated from experiments. Several $T=0$ predictions for ultracold Fermi gases were made in the past by means of quantum Monte Carlo (QMC) techniques Carlson et al. (2003); Chang et al. (2004); Carlson and Reddy (2005, 2008); Astrakharchik et al. (2004); Gezerlis et al. (2009); Carlson et al. (2011); Braun et al. (2015), and were later confirmed by experiments for both the equal-mass Ku et al. (2012) and unequal-mass Kohstall et al. (2012) cases. The variational Monte Carlo (VMC) and the diffusion Monte Carlo (DMC) methods allow one to accurately solve for the ground state of strongly interacting many body systems For the case of the central interactions in this work the diffusion Monte Carlo approach is exactly equivalent to, and occasionally referred to as . The QMC family of approaches has been widely used in the study of diverse physical systems such as condensed matter, quantum chemistry, nuclear physics, and cold atomic physics Ceperley (1995); Foulkes et al. (2001); Kolorenč and Mitas (2011); Roggero et al. (2013); Carlson et al. (2015); Schonenberg and Conduit (2017); Gandolfi et al. (2018); Curry et al. (2023). Previous studies explored the properties of strongly interacting fermions at unitarity Gandolfi et al. (2011) however, the theoretical study of the unequal mass system has not received effort similar to the equal-mass case.
In this paper we focus on a Fermi-Fermi mixture with a mass ratio $m_{h}/m_{l}$=6.5 corresponding to the 6Li-40K mixture, and we study the BCS-BEC crossover by employing QMC techniques; in particular in this work we compute the properties of the unpolarized system with an equal number of heavy and light atoms. We first compute the equation of state that is crucial to fit the density functionals used to compute the properties of trapped larger systems Forbes et al. (2011); Bausmerth et al. (2009); Forbes et al. (2012). The preliminary results of these calculations were presented in Gandolfi (2014), although we have considerably expanded upon that early work, addressing a number of physical observables. We compute the energy per particle as a function of the scattering length of the two-body interaction by paying particular attention to the effect of the finite effective range of the potential. We also compute the pair-distribution function between heavy and light particles as well as the momentum distribution. These quantities are useful for extracting the contact parameter that can be measured in experiments, similar to the equal-mass case Tan (2008a, b, c). We then simulate the system with very small polarization in order to study the quasiparticle dispersion as a function of the interaction. These calculations serve to carefully explore the properties of the 6Li-40K system and provide stringent theoretical predictions for future experiments.
II Model
We model our system by considering point-like interacting particles with the following Hamiltonian:
$H=\frac{-\hbar^{2}}{2m_{l}}\sum_{i=1}^{N_{l}}\nabla_{i}^{2}+\frac{-\hbar^{2}}{2m_{h}}\sum_{j^{\prime}=1}^{N_{h}}\nabla_{j^{\prime}}^{2}+\sum_{i<j^{\prime}}v(r_{ij^{\prime}})\,,$ (1)
where the sum $i$ is over $N_{l}$ light particles with mass $m_{l}$ and the $j^{\prime}$ sum is over $N_{h}$ heavy particles with mass $m_{h}$. $v(r_{ij^{\prime}})$ is an $s-$wave short-range interaction acting only between fermions of different species. The form of $v(r)$ we use is the Pöschl-Teller already employed in several previous QMC calculations Carlson et al. (2003); Chang et al. (2004); Carlson and Reddy (2005); Gezerlis et al. (2009); Morris et al. (2010); Gandolfi et al. (2011); Forbes et al. (2011, 2012):
$v(r)=-v_{0}\frac{\hbar^{2}}{m_{r}}\frac{\mu^{2}}{\cosh^{2}(\mu r)}\,,$ (2)
where $m_{r}$ is the reduced mass, and the parameters $\mu$ and $v_{0}$ are tuned in order to reproduce the effective range $r_{e}$ and the scattering length $a$ of the interaction. For this potential, the unitary limit corresponding to the zero-energy ground-state between two particles is achieved with $v_{0}=1$ and $r_{e}=2/\mu$. We consider a system in the dilute limit, meaning that the interparticle distance $r_{0}\gg r_{e}$, where $r_{0}=(9\pi/4)^{1/3}/k_{F}$ and $k_{F}$ is the Fermi momentum of the system. Most of the results presented in this work were obtained by fixing $r_{e}k_{F}\approx 0.03$, but in several cases we considered different values to check effects due to the effective range of the potential.
The ground state of the system is solved through the use of QMC techniques, in particular VMC and DMC. The many-body wave function  Carlson et al. (2003); Chang et al. (2004) is given by
$\Psi_{T}=\left[\prod_{i<j^{\prime}}f(r_{ij^{\prime}})\right]\Phi_{\rm BCS}\,.$ (3)
The Jastrow function $f(r)$ acts between particles with different masses, and it is obtained by solving the equation
$-\frac{\hbar^{2}}{2m_{r}}\nabla^{2}f(r)+v(r)f(r)=\lambda f(r)\,.$ (4)
The parameter $\lambda$ is obtained by imposing the boundary condition $f(r>d_{0})=1$, where the healing distance $d_{0}$ is a variational parameter. The correct boundary conditions of the wave function are guaranteed by constraining $d_{0}\leq L/2$, where $L$ is the size of the simulation box, but we find that a typical good choice is $d_{0}\approx L/10$. By solving Eq. (4) we also ensure the correct behavior of the wave function at small distances so that the Jastrow function $f$ is defined to have $\partial f/\partial r=0$ at the origin. This condition must be satisfied in order to have a finite kinetic energy in the origin when computing $\nabla^{2}f=f^{\prime\prime}(r)+2f^{\prime}(r)/r$. This condition is necessary to avoid spurious contributions to the energy computed using QMC.
The antisymmetric part $\Phi_{\rm BCS}$ is a particle-projected BCS wave function including pairing correlations (see the Appendix of Ref. Gandolfi et al. (2009)). It is given by
$\Phi_{\rm BCS}={{\cal A}}[\phi(r_{11^{\prime}})\phi(r_{22^{\prime}})...\phi(r_{N_{l}N_{h}})]\,.$ (5)
The operator ${\cal A}$ antisymmetrizes like particles, the unprimed coordinates are for light particles, the primed ones are for heavy particles, and $N_{l}=N_{h}=N/2$ for the unpolarized case. The pairing function is expressed as
$\phi({\bm{r}})$ $=$ $\beta(r)+\sum_{\bm{n},I\leq I_{C}}a(k_{\bm{n}}^{2})\exp[i\bm{k}_{\bm{n}}\cdot\bm{r}]~{},$
$\beta(r)$ $=$ $\tilde{\beta}(r)+\tilde{\beta}(L-r)-2\tilde{\beta}(L/2)~{},$
$\tilde{\beta}(r)$ $=$ $[1+\gamma br]\ [1-\exp(-dbr)]\frac{\exp(-br)}{dbr}~{}.$ (6)
The function $\beta(r)$ has a range of $L/2$, and the value of $\gamma$ is chosen such that $\beta(r)$ has zero slope at the origin. The sum in the pairing function $\phi(\bm{r})$ is over $\bm{k}_{\bm{n}}=\frac{2\pi}{L}(n_{x},n_{y},n_{z})$, which imposes a shell structure according to $I=n_{x}^{2}+n_{y}^{2}+n_{z}^{2}$ that we cutoff at some value $I_{C}$. Note that if $\beta(r)=0$, and $a(k_{n}^{2})=0$ for $|\bm{k}_{n}|>k_{F}$, $\Phi_{\rm BCS}=\Phi_{FG}$, where the latter is a Slater determinant wave function describing the noninteracting Fermi gas in the normal phase. As $\Phi_{FG}$ and $\Phi_{\rm BCS}$ are orthogonal, we can study the Fermi gases both in the superfluid and in the normal phase Lobo et al. (2006); Pilati and Giorgini (2008). The pairing wave function used here contains 12 free parameters that have been optimized by minimizing the energy of the system using VMC and following the strategy of Ref. Sorella (2001). In the case of strong interactions, when $1/ak_{F}\geq 2$, we found that using the pairing function as in Ref. Astrakharchik et al. (2004), gives lower energies. Practically this involves finding the solution to the two-body Schrödinger equation following a procedure similar to what was discussed for Eq. (4), except in this case we substitute $\lambda$ with the two-body binding energy. Instead, in the BCS case for $11/ak_{F}\leq-$ 1, the BCS wave function $\Phi_{BCS}$ gives almost the same energy as $\Phi_{FG}$ as the pairing becomes less and less important.
Polarized systems can be simulated by extending the wave function to include single-particle states for the unpaired particles Carlson et al. (2003); Chang et al. (2004). Note that the wave function as previously described cannot reproduce more exotic phases like the LOFF phase or $p$-wave pairing.
The ground state of the system is then solved by projecting the wave function in imaginary time by means of the DMC technique (see, for example, Ref. Chang et al. (2004)). It is important to note that the DMC algorithm strictly provides an upper bound to the energy, and thus it is a variational calculation within the fixed-node approximation used to control the fermion sign problem. However, if the global structure of the wave function is carefully optimized, the DMC provides a very accurate estimate of the ground state. The accuracy of the DMC has been tested by comparing the results of the equal-mass unitary Fermi gas with the auxiliary field quantum Monte Carlo (AFQMC) Carlson et al. (2011). For purely attractive interactions, in AFQMC the unpolarized system is solved exactly with no sign problem, although AFQMC calculations do need to account for lattice discretization effects Curry et al. (2024). The agreement between DMC and AFQMC is within 4%-5% giving us the confidence that once the wave function is carefully optimized, the upper bound to the energy given by DMC will be very close to the energy of the ground state. It is straightforward to extend the DMC to deal with systems with unequal masses or with polarization by modifying the wave function, and the systematic error given by the fixed-node approximation should provide the same accuracy as for the unpolarized case if the trial wave function correctly describes the phase of the system. Note that for polarized systems or for unpolarized systems with unequal masses the AFQMC suffers from a severe sign problem, and the constrained path used to control it does not provide an upper bound to the energy as the fixed-node approximation does for DMC. It should also be noted that all of the uncertainties presented on the calculations in this work are statistical in nature, arising from observables in QMC being computed as averages over many imaginary-time propagated copies of the system under study.
III BCS-BEC crossover
We simulate a system of 66 fermions, as this has been shown to provide a very accurate description of fermionic systems in the thermodynamic limit Forbes et al. (2011); Gandolfi et al. (2011); Palkanoglou et al. (2020). In addition it has been shown Forbes et al. (2012) that near unitarity finite-size effects are minimal once the number of particles is $\geq 40$. However, since we are investigating the full crossover region, the choice of 66 particles also has the benefit of minimizing any shell effects in the weakly interaction regime, as 66 particles are a closed-shell configuration in our wavefunction. A final benefit of choosing this particular number is that results found using a BCS-like wave function can be directly compared to the system in the normal phase. In the $\Phi_{FG}$ wave function the single-particle orbitals are plane waves, and the energy of 33 free fermions is very close to the infinite limit.
Figure 1: (color online) The BCS-BEC crossover for the unpolarized 6Li-40K mixture as a function of $1/a\,k_{F}$. QMC results are shown for two potentials with different effective ranges (blue circles for $\mu/k_{F}=60$ and green squares for $\mu/k_{F}=24$). The energy is in units of the Fermi gas energy $E_{FG}$ after the subtraction of the two-body binding energy per particle on the BEC side. The orange dot-dashed line shows the perturbation expansion of a weakly attractive Fermi gas from Eq. (7)
, and the red dashed line shows the equation of state of a repulsive Bose gas from Eq. (8), with a dimer-dimer scattering length $a_{dd}=0.886(4)a$.
We show in Fig. 1 the equation of state of the unpolarized 6Li-40K mixture as a function of $1/ak_{F}$. For negative values of $a\,k_{F}$ the two-body interaction does not admit a two-body bound state, and for positive values the two-body binding energy per particle is subtracted from the total energy of the system. The results are in units of the non-interacting Fermi gas, $E_{FG}=3E_{F}/5$, where $E_{F}=\frac{\hbar^{2}}{4m_{r}}k_{F}^{2}$. It is clear that on the BCS side the energy per particle approaches the limit of the noninteracting Fermi gas, which can additionally be seen by comparison with the perturbative expansion of a weakly interacting Fermi gas Huang and Yang (1957) given by,
$\frac{E}{E_{FG}}=1+\frac{10}{9\pi}ak_{F}+\frac{4}{21\pi^{2}}(11-2\ln 2)(ak_{F})^{2}+\cdots.$ (7)
On the BEC side, the energy per particle quickly becomes comparable to the two-particle binding energy showing the formation of strongly bound molecules, and $(E-E_{2}/2)/E_{FG}$ rapidly tends to zero. All the points have been computed by setting the effective range of interaction corresponding to $\mu/k_{F}=60$ and tuning $v_{0}$ to have the desired scattering length (blue circles in Fig. 1). For many points we investigated the effect of a finite effective range and repeated the calculation using $\mu/k_{F}=24$ (green squares). We note that the results become independent of the effective range for $1/|a\,k_{F}|\geq 1$, as seen by the essentially constant energy in the deep BEC regime.
As discussed above for $11/ak_{F}>$ 1, we found that the best variational wave function is provided by assuming the pairing wave function is the two-body solution as in Ref. Astrakharchik et al. (2004). By analyzing the structure of the wave function we can conclude that for $1/ak_{F}\geq 2$ the system made of attractive fermions is very well modeled by a weakly repulsive Bose gas. In this regime, if we use the pairing wave function in Eq. (II), the energy of the system is higher, and we would likely need many more plane waves to correctly describe the short-range behavior of the two-particle solution.
Figure 2: (color online) The BCS-BEC crossover for the unpolarized 6Li-40K mixture as a function of $1/a\,k_{F}$ near the unitary limit. Our results are compared to a fit given by Eq. (9) as discussed in the text.
The energy of a repulsive Bose gas, in which bosons are made by two fermions with different masses, can be well approximated by a mean-field expansion Lee et al. (1957),which for unequal masses is given by
$\frac{E/N-E_{2}/2}{E_{FG}}$ $=$ $\frac{10}{9\pi}a_{dd}k_{F}\frac{m_{h}m_{l}}{(m_{h}+m_{l})^{2}}$ (8)
×$\times\left[1+\frac{128}{15\sqrt{6\pi^{3}}}(a_{dd}k_{F})^{3/2}+\dots\right],$
where $a_{dd}$ is the boson-boson scattering length, whose value is obtained by fitting to our QMC calculations. The results in the strongly interacting regime are shown in the right inset of Fig. 1 for the mass ratio considered. The fit is made in the $1/ak_{F}\geq 2$ region, giving $a_{dd}/a=0.886(4)$, in good agreement with the few-body analysis of Refs. Petrov et al. (2005); Levinsen and Petrov (2011). The agreement between the QMC results and the mean field expansion of Eq. (8) starts to fail for $21/ak_{F}<$ 2, indicating that the system is no longer similar to a Bose gas, as also suggested by looking at the structure of the many-body wave function.
In Fig. 2 we show the crossover around the unitary limit. Close to unitarity, the equation of state can be expanded as Tan (2008b)
$\frac{E}{E_{FG}}=\xi-\frac{\zeta}{k_{F}a}-\frac{5\nu}{3(k_{F}a)^{2}}+\dots\,.$ (9)
Figure 3: (color online) The energy of the 6Li-40K mixture as a function of the effective range $r_{e}$ of the interaction in the unitary limit. The energy is in units of the Fermi gas energy $E_{FG}$. The slope of the extrapolation is similar to what was found in the equal-mass case with lattice calculations Carlson et al. (2011). The details of the fit to the QMC results are discussed in the text.
The fit to our QMC results near unitarity gives ${\xi=0.3726(6)}$, ${\zeta=0.900(2)}$, and ${\nu=0.46(1)}$. Our result for $\xi$ differs appreciably from the equal-mass case near unitarity ($\xi=0.383(1)$ from Gandolfi et al. (2011)), which is expected to be a result of non-zero total momentum pairs in the ground state Gezerlis et al. (2009). However, the values of $\zeta$ and $\nu$ are very similar to what was found for the equal-mass case Ref. Gandolfi et al. (2011). Our results are in good qualitative agreement with the calculation of Ref. Diener and Randeria (2010), deviating from mean-field calculations which predict no difference between the equal- and unequal-mass cases.
The various tests we performed indicate that around ${1/a\,k_{F}\approx 0}$ the extrapolation to $r_{e}\rightarrow 0$ just shifts the energy and does not change the curvature of the equation of state in the region where $11/|a\,k_{F}|<$ 1. There, only the $\xi$ parameter is particularly sensitive to $r_{e}$, which is also the case for the equal-mass system Gandolfi et al. (2011). It is interesting to compare the finite-range extrapolation for the unequal-mass mixture to the equal-mass case. The comparison is warranted for the 6Li-40K system because the resonance is narrower than for equal masses Naik et al. (2011), and consequently, the unitary limit could be “less universal”, and the effect of the effective range could be more important. We performed several calculations at unitarity for different values of $r_{e}$ in order to test the finite-range effects and extrapolated to zero effective range as shown in Fig. 3. The QMC results were fit using the linear function
$\frac{E}{E_{FG}}=\xi+c\,r_{e}k_{F}\,,$ (10)
and the best estimate given by QMC is $\xi=0.368(1)$. Similar to the above case near unitarity, our value for $\xi$ at the unitarity limit is slightly lower than what is found in equal-mass calculations carried out on the lattice ($\xi=0.372(3)$ from Carlson et al. (2011). This is further evidence of the contribution of nonzero total momentum pairs in the ground state, as found in Gezerlis et al. (2009). The value of $c=0.13(1)$ gives the slope of the extrapolation, and it is rather intriguing that the slope of the energy as a function of the effective range is very similar to the equal-mass case Carlson et al. (2011). This fact, along with the similarities to higher order terms from Eq. (9), should be explored in future investigations into mass-imbalanced cold atomic systems.
As discussed above, there are many reasons to expect different physics between the unequal-mass Fermi gas that we study and the more common equal-mass case. At the level of the equation of state the two systems display certain similarities (higher-order terms in the equation of state), as well as differences (the Bertsch parameter $\xi$ is quantitatively different near and at the unitary limit). To investigate further similarities and differences between the equal- and unequal-mass Fermi gases, we consider a number of different physical observables.
Figure 4: (color online) Heavy-light pair-distribution function for different two-body interactions indicated in the legend. From left to right they correspond to $41/ak_{F}=-$ 4, $-1$, $-0.5$, $-0.1$, 00, $0.1$, and $0.5$. In the inset we show the same function multiplied by $(k_{F}r)^{2}$ for selected interaction strengths corresponding to $41/ak_{F}=-$ 4, $-1$, $-0.5$, 00, and $0.5$ (from the bottom to the top).
We first consider the pair-distribution function $g(r)$ between heavy and light particles, as shown in Fig. 4. $g(r)$ is typically calculated by a mixed estimate of the form
$g(r)=\sum_{i<j^{\prime}}\frac{\left\langle\Psi_{0}\right|\delta(r_{ij^{\prime}}-r)O_{ij^{\prime}}^{P}\left|\Psi_{T}\right\rangle}{\left\langle\Psi_{0}|\Psi_{T}\right\rangle},$ (11)
where $\Psi_{0}$ is the DMC projected ground state and, for our purposes ($s$-wave central interaction) the $O_{ij^{\prime}}^{p}$ operator is equal to 1 Gezerlis and Carlson (2010). The sum is only over pairs of particles with different masses.
The function $g(r)$ exhibits a large peak at short distances, induced by the strong attractive interaction in the $s$-wave channel. In the weak-coupling limit the value of the peak is smaller as expected, and is clear in our results. In the inset of Fig. 4 we show $(k_{F}r)^{2}g(r)$, which can be used to extract the contact parameter, as shown in Ref. Gandolfi et al. (2011). The limit of $(k_{F}r)^{2}g(r)$ for short distances approaches zero in the BCS limit as the short-range correlations are weaker when the strength of the attractive interaction decreases. On the other hand, in the strong-coupling regime where $0.51/ak_{F}=$ 0.5, the limit of $(k_{F}r)^{2}g(r)$ at small distances rapidly increases. Our results for unitarity (red diamonds in Fig. 4) can be compared with the equal-mass case shown in Fig. 4 of Gandolfi et al. (2011). One notable difference is that in the equal-mass case $g(r)$ decays with increasing $k_{F}r$, and we find this effect to be dramatically reduced for the unequal-mass case. By looking at the inset of Fig. 4, at unitarity we see that $(k_{F}r)^{2}g(r)$ is essentially constant over the considered range of $k_{F}r$. This can be interpreted as the heavy-light pairs remaining farther apart from one another, potentially due to the higher inertia of the heavy particles, compared to the equal-mass case.
Figure 5: (color online) Momentum distribution $n(k)$ for the unequal-mass system at unitarity. We calculate $n(k)$ using only the light particles (solid line) and using only the heavy particles (dashed line). In the inset we plot $(k/k_{F})^{4}n(k)$ to show the tail behavior of the momentum distribution. The dotted black line shows the contact parameter $C$ derived from our calculated equation of state around the unitary limit, following the prescription from Werner and Castin (2012)
Figure 6: (color online) Momentum distribution $n(k)$ for different interaction strengths (same color convention as in Fig. 4), using only the light particles. In the inset we plot $(k/k_{F})^{4}n(k)$ to show the tail behavior of the momentum distribution. The results for unitarity (red diamonds) also appear in Fig. 5.
In Figs. 5 and 6 we show momentum distributions for the mass imbalanced system, calculated as the Fourier transform of the one-body density matrix by,
[…]
where the integral over the solid angle is handled stochastically and the integral over $\delta r=|\bm{r}^{\prime}_{n}-\bm{r}_{n}|$ is carried out on a line to avoid statistical errors Gezerlis and Carlson (2010). The primed coordinate refers to a single particle that is moved along this line in order to carry out the integration over $d\delta r$. Equation (12) can then be computed for each particle in our system, and an average value can be taken. In Fig. 5 we first investigate the momentum distribution for the system at unitarity, averaging over only the light or heavy particles independently. We also investigate the tail behavior of the momentum distribution by plotting $(k/k_{F})^{4}n(k)$. We find in both cases a value for Tan’s contact parameter that agrees with the expected result from Werner and Castin (2012), given our equation of state in Fig. 1. Figure 5 also highlights a subtle difference between the momentum distributions of the heavy and light particles, with the heavy-particle calculations showing a noticeable decrease for low-momentum values. This feature is absent in studies of the equal-mass Fermi gas.
Figure 7: (color online) Quasiparticle dispersion at different scattering lengths. The upper and lower curves show results for light (solid lines) and heavy (dashed lines) quasiparticles as a function of the momentum. We show results for the unitary limit (red diamonds) as well as the BCS region ($ak_{F}<0$) and the strong interacting BEC region ($ak_{F}>0)$. All results correspond to a mass ratio $m_{h}/m_{l}=6.5$.
Since considering either the light or heavy particles alone gives the same value for the contact, we can investigate the crossover behavior of the momentum distribution, as shown in Fig. 6, considering only the light particles. In the very weak coupling region with $41/ak_{F}=-$ 4, $n(k)$ is very similar to the non interacting Fermi gas momentum distribution, where $n(k)=1$ for $k<k_{F}$ and zero otherwise. Moving to strong couplings the tail of $n(k)$ is longer, as expected, due to the strong pairing effects. This is clearly evident in the inset of Fig. 6 where the function $(k/k_{F})^{4}n(k)$ is shown. The inset clearly shows the $C/k^{4}$ universal behavior predicted by Tan Tan (2008a, b, c), where $C$ is the contact, regardless of interaction strength. The value of $C$ we find from the tail of $n(k)$ agrees with the value that can be computed from a derivative of the equation of state with respect to the scattering length Werner and Castin (2012). It is also interesting to note that the value of $C$ we find for the unequal-mass case agrees within error with what is found for the equal-mass case Gandolfi et al. (2011). This highlights a fundamental similarity between the equal- and unequal-mass Fermi gases.
Last, the quasiparticle dispersion is computed by adding an unpaired particle to the system Carlson et al. (2003); Chang et al. (2004); Gezerlis et al. (2009); Gandolfi et al. (2008); Gezerlis and Carlson (2010). We carried out simulations of 67 particles where the unpaired particle is characterized by a finite momentum $|k|$ and mass $m_{l}$ or $m_{h}$. We then combine the result with the energy of the fully paired unpolarized system of 66 particles and define the quasiparticle energy as
$\epsilon(k)/E_{F}=[E_{k}(67)-E(66)]/E_{F}\,,$ (13)
where $k$ is the momentum of the extra particle, and $E(67)$ and $E(66)$ are the total energies of the systems with 67 and 66 particles. The two simulations were performed at the same total density. In Ref. Gezerlis et al. (2009) the quasiparticle dispersion was computed by comparing the energy of the two systems at the same background density, then at constant volume. Here we have fixed the total density in order to avoid small effects due to the effective range of the interaction. However, the finite-size effects should be very small as the volume of the simulation box is quite large.
The results of the quasiparticle dispersion in Fig. 7 show that the dispersions of light and heavy particles are very different from what was already found at unitarity Gezerlis et al. (2009). For each case the quasiparticle dispersion is lower on the BCS side and increases by crossing unitarity and moving to the BEC side. In particular for light quasiparticles the position of the minimum changes from about $(k/k_{F})^{2}=0.5$ at $ak_{F}=-2$ to $(k/k_{F})^{2}=0$ at $ak_{F}=2$. Our results at unitarity agree with previous findings Gezerlis et al. (2009) that the quasiparticle dispersion for the unequal-mass Fermi gas differs dramatically from the equal-mass case Carlson and Reddy (2005). We have also shown that the differences in the quasiparticle dispersions for heavy and light particles are seen throughout the BCS-BEC crossover region.
IV Conclusions
In this paper we showed QMC results exploring the BCS-BEC crossover of the mass-imbalanced 6Li-40K Fermi mixture. QMC techniques were used in the past to give the best upper bound of the energy of strongly interacting fermions in the unitary limit. We made use of QMC to compute the equation of state as a function of the two-body scattering length, studying the crossover from the weakly interacting limit where the system is well described by the BCS theory to the strong coupling limit where fermions form molecules, and the system is well described by a weakly repulsive Bose gas. We carefully checked finite-range effects close to the unitary limit and provided an accurate fit of the equation of state based on our QMC calculations. We also showed other properties of the unpolarized superfluid system, namely, the pair-distribution function and the momentum distribution. Finally, by considering a system with very small polarizations, we computed the heavy- and light-quasiparticle dispersions for different strengths of the interaction. Throughout we also paid careful attention to the differences and similarities between the equal- and unequal-mass Fermi gases, finding several important results. Our calculations of the equation of state revealed that the higher order terms in the equation of state are very similar. We also found that the universal contact parameter is identical to that in the equal-mass case. However, we did find several key differences from the equal-mass case, such as the Bertsch parameter, quasiparticle dispersion, and tail of the pair-distribution function. Our results should provide a comprehensive framework for future theoretical and experimental investigations into mass-imbalanced Fermi gases across the full range of the BCS-BCS crossover.
Acknowledgments
The work of S.G. is supported by the U.S. Department of Energy, Office of Nuclear Physics, under Contract No. DE-AC52-06NA25396, and by the Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) NUCLEI program. The work of R.C. and A.G. was supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada and the Canada Foundation for Innovation (CFI). R.C. was also supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under Project No. 20220541ECR. Computational resources were provided by the Los Alamos National Laboratory Institutional Computing Program, which is supported by the U.S. Department of Energy National Nuclear Security Administration under Contract No. 89233218CNA000001. Computer time was also made available by the National Energy Research Scientific Computing Center (NERSC) and SHARCNET through the Digital Research Alliance of Canada.